A manufacturer estimates that when q units of a certain commodity are produced, the total cost will be



C(x) rand where



C(q) = q2



25 + 80 000 − 104q.



Answer the following questions:



(i) Use marginal analysis to determine the production level at which the cost will be a minimum.



(ii) Determine the minimum cost.


1
Expert's answer
2022-04-03T14:45:20-0400

"C(q)=25q^2+80000-104q"

"C'(q)=50q-104=0"

"q=104\/50=2.08"

This is production level at which the cost will be a minimum.

The minimum cost is

"C(2.08)=25\times2.08^2+80000-104\times2.08=108.16+80000+216.32=80324.5"